### v3

@@ 4.1 Data Description
Table 1: Sample distribution information of Datasets
@@ 16 Stone-Steel-Towers 93 - - - -
4.2 Experiment Design
The experimental setup encompasses two comparative segments for classification evaluation. The first part involves basic model comparison experiments, while the second focuses on comparing various tuning strategies.
In the basic model comparison part, we partition each target HSI dataset into two subsets: the training set and test set. We randomly extract 150 samples from each category as test samples and take the rest as the training samples in the Pavia University dataset and Houston University. For the Indian Pines dataset, we sample a small number of pixels from classes with scarce samples to ensure a more balanced representation for the experiments. Concretely, we randomly select 10 samples from each of the following six classes: Alfalfa (class 1), Grass/trees (class 5), Grass/pasture-mowed (class 7), Oats (class 9), Buildings-grass-trees (class 15), and Stone-steel towers (class 16). For the remaining classes, we sample 150 pixels from each class. In this Section  4.4, we compare our proposed Tri-Former model with other state-of-the-art algorithms.
In addition, we compare the proposed AGT approach with other popular tuning methods on HSIs collected from different sensors. The tuning strategies comparison results are presented in Section 4.6 and Section 4.7. These experiments involve data from two source sensors: Salinas (captured by AVIRIS) and Pavia Center (captured by ROSIS). For the model trained on the data collected by one sensor, we fine-tune it on the target data collected by another sensor with a limited number of samples. We design different experimental groups, each containing 25, 50, and 75 labeled pixels per class, respectively. Furthermore, in Section 4.8, we evaluate the proposed method by cross-modality fine tuning experiments. Specifically, the basic Tri-Former model pretrained on RGB images is fintuned on different HSI datasets with different tuning method.
@@ 4.2 Experiment Design
4.3 Implementation Details and Evaluation Metrics
Implementation details: The optimization of the proposed method consists of two stages. The entire learning process is conducted on the server with 4 NVIDIA 3090 GPUs. During training, the training samples are cropped as 27x27 spatial resolution patches, with a batch size of 96 on all three training dataset. We utilize the AdamW optimizer with a learning rate and weight decay value of $1\times 10^{-3}$ and $1\times 10^{-5}$, respectively. The learning rate decays from $1\times 10^{-3}$ to $1\times 10^{-6}$ following the cosine scheduler. The warming-up stage comprises 5 epochs and all the models are trained for 300 epochs. To further tune the final network, we crop patches with spatial resolutions of 13x13 and 27x27, which are fed into the auxiliary branch and basic model, respectively. Batch sizes are set to 12 for all datasets during tuning.
4.4 Comparison between Tri-Former and state-of-the-art methods
In this section, we conduct a comparative analysis of our proposed Tri-Former against five ConvNet-based and transformer-based HSI classification methods. The comparisons are based on official implementations of these methods: 3D-CNN [^2] https://github.com/eecn/Hyperspectral-Classification Chen et al. (2016), LWNet [^3] https://github.com/hkzhang91/LWNet Zhang et al. (2019), SpectralFormer [^4] https://github.com/danfenghong/IEEE_TGRS_SpectralFormer Hong et al. (2021), GraphGST [^5] https://github.com/yuanchaosu/TGRS-graphGST Jiang et al. (2024) and SSFTT [^6] https://github.com/zgr6010/HSI_SSFTT Sun et al. (2022). The experimental results are presented in Table 2 to Table 4, and the visual comparisons are shown in Figure 8 to Figure10.
@@ 4.4 Comparison between Tri-Former and state-of-the-art methods
Figure 8: Results on Indian Pines dataset. (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) Tri-Former.
@@
Table 2: Comparison Experimental Results on Indian Pines Using 150 Training Samples Each Class
@@
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT Tri-Former
1 91.67 97.22 77.78 100.00 100.00 97.14
2 94.84 96.01 90.06 94.91 94.99 99.45
3 99.12 98.82 94.26 99.26 99.12 99.71
4 100.00 100.00 98.85 74.89 100.00 100.00
5 100.00 100.00 99.10 96.99 100.00 91.45
6 100.00 99.83 98.97 97.76 99.83 100.00
7 100.00 100.00 94.44 94.44 100.00 100.00
8 100.00 100.00 100.00 100.00 100.00 100.00
9 100.00 100.00 100.00 100.00 100.00 100.00
10 99.39 100.00 96.35 99.15 99.51 99.51
11 92.41 98.39 90.54 93.02 95.05 98.13
12 98.42 99.77 96.61 96.61 99.55 99.56
13 100.00 100.00 100.00 97.95 100.00 100.00
14 98.92 99.91 98.92 99.46 98.82 100.00
15 66.49 70.74 68.35 100.00 89.63 91.76
16 90.36 91.57 84.33 97.59 83.13 89.16
OA 95.23 97.46 93.08 96.02 97.12 98.41
AA 95.73 97.02 93.04 96.38 97.54 97.87
K 94.47 97.05 91.97 95.41 96.66 98.17
@@
Table 3: Comparison Experimental Results on Pavia University Using 150 Training Samples Each Class
@@
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT Tri-Former
1 91.09 96.06 83.13 98.06 94.69 99.98
2 92.05 95.69 92.83 98.03 98.74 99.91
3 92.45 95.95 89.34 92.82 98.50 100.00
4 98.55 98.68 97.60 99.11 95.05 99.18
5 99.52 99.92 100.00 100.00 97.43 100.00
6 93.14 98.09 90.32 90.92 99.96 100.00
7 97.48 99.76 96.99 100.00 100.00 99.92
8 95.20 99.11 92.63 95.95 99.08 99.77
9 97.28 98.82 99.65 99.74 99.88 100.00
OA 93.26 96.86 91.65 98.06 98.03 99.88
AA 95.20 98.01 93.61 98.18 98.15 99.86
K 91.13 95.84 89.00 97.41 97.38 99.83
The quantitative analysis results of the aforementioned approaches on the Indian Pines dataset are listed in Table 2. Table 2 shows that the proposed Tri-Former achieves the highest OA of 98.41%, outperforming other ConvNet-based HSI classification methods like 3D-CNN (95.23%), LWNet (97.46%), SpectralFormer (93.08%), GraphGST (96.02%) and SSFTT (97.12%). Among them, SSFTT incorporation of convolutional operations within the transformer structure, which benefits its performance. This improvement further substantiates some of the points we have mentioned in this work. Introducing convolutional operations to construct hybrid transformer models for processing HSI is a more suitable approach. The proposed Tri-Former further improved the OA metric of SSFTT by 1.29%, demonstrating its remarkable capability to accurately classify HSIs in this challenging dataset.
Figure 8 shows the corresponding visual results. At the top, a color-composite image is presented, accompanied by color codes that represent different types of ground objects. The bottom-right corner features a magnified view, aimed at providing a clearer comparison of the classification performance of different methods. From Figure 8, we can see that all methods exhibit varying degrees of misclassification. Our proposed Tri-Former demonstrates the least amount of misclassification, approaching the ground truth closely.
@@
Figure 9: Results on Paiva University. (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) Tri-Former.
Similar trends are also demonstrated in the experiments conducted on the Pavia University and Houston University datasets. When evaluated on the Pavia University dataset, the proposed Tri-Former algorithm surpasses the state-of-the-art method SSFTT, achieving remarkable improvements in key evaluation metrics. Specifically, we observe substantial enhancements in Overall Accuracy (OA), Average Accuracy (AA), and Kappa (K) coefficient by 1.85%, 1.71%, and 2.45%, respectively. On the Houston University dataset, where the potential for performance improvement is considerably limited, our Tri-Former still manages to achieve the best performance among all methods compared. Figure 9 and Figure 10 are the corresponding visual comparison results. On these two datasets, especially on the Pavia University dataset, the proposed Tri-Former produces classification result maps that closely resemble the groundtruth.
Figure 10: Results on Houston University. (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) Tri-Former.
@@
Table 4: Comparison Experimental Results on Houston University Using 150 Training Samples Each Class
@@
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT Tri-Former
1 96.64 98.64 96.00 78.47 98.00 99.64
2 95.20 99.01 89.31 83.51 97.10 97.92
3 99.82 100.00 99.09 84.82 99.09 100.00
4 97.62 99.54 86.47 79.43 97.90 99.73
5 99.45 100.00 95.88 97.43 100.00 99.91
6 100.00 100.00 86.86 88.57 100.00 100.00
7 91.68 96.06 73.23 62.88 98.66 99.11
8 94.70 96.71 85.56 54.39 97.62 97.17
9 96.82 97.19 85.57 54.38 99.82 98.82
10 98.33 99.44 91.55 70.10 100.00 99.63
11 97.05 100.00 92.35 51.34 100.00 97.05
12 96.12 98.25 86.98 60.85 96.21 100.00
13 95.61 100.00 82.45 86.20 99.37 100.00
14 100.00 100.00 100.00 94.96 100.00 100.00
15 100.00 100.00 97.65 86.67 100.00 100.00
OA 96.58 98.70 89.19 94.42 98.69 99.05
AA 97.14 98.99 89.93 95.13 98.92 99.26
K 96.29 98.59 88.28 93.94 98.57 98.97
4.5 Inference speed analysis on Tri-Former

### v4

@@ IV-A Data Description
TABLE I: Sample distribution information of Indian Pines & Pavia University Houston & University Datasets
@@ 16 Stone-Steel-Towers 93 - - - -
TABLE II: Sample distribution information of WHU-Hi Datasets
WHU-Hi-LongKou WHU-Hi-HanChuan WHU-Hi-HongHu
1 Corn 34511 Strawberry 44735 Red roof 14041
2 Cotton 8374 Cowpea 22753 Road 3512
3 Sesame 3031 Soybean 10287 Bare soil 21821
4 Broad-leaf soybean 5353 Sorghum 3064 Cotton 163285
5 Narrow-leaf soybean 4151 Water spinach 1200 Cotton firewood 6218
6 Rice 11854 Watermelon 4533 Rape 44557
7 Water 67056 Greens 5903 Chinese cabbage 24103
8 Roads and houses 7124 Trees 17978 Pakchoi 4054
9 Mixed weed 5229 Grass 9469 Cabbage 10819
10 - - Red roof 10516 Tuber mustard 12394
11 - - Gray roof 16911 Brassica parachinensis 11015
12 - - Plastic 3679 Brassica chinensis 8954
13 - - Bare soil 9116 Small Brassica chinensis 22507
14 - - Road 18560 Lactuca sativa 7356
15 - - Bright object 1136 Celtuce 1002
16 - - Water 75401 Film covered lettuce 7262
17 - - - - Romaine lettuce 3010
18 - - - - Carrot 3217
19 - - - - White radish 8712
20 - - - - Garlic sprout 3486
21 - - - - Broad bean 1328
22 - - - - Tree 4040
Total 146683 Total 255241 Total 381693
@@ 9 Mixed weed 5229 Grass 9469 Cabbage 10819
WHU-Hi-LongKou dataset [51] was captured on July 17, 2018, in Longkou Town, Hubei province, China. The imaging was conducted using a DJI Matrice 600 Pro UAV, equipped with an 8-mm focal length Headwall Nano-Hyperspec sensor, flying at an altitude of 500 meters. The resulting dataset consists of images with dimensions of 550 × 400 pixels and a spatial resolution of about 0.463 meters. It spans 270 spectral bands, covering wavelengths from 400 to 1000 nm. The scene mainly includes agricultural fields with six types of crops: corn, cotton, sesame, broad-leaf soybean, narrow-leaf soybean, and rice.
WHU-Hi-HanChuan dataset [51] was gathered on June 17, 2016 in Hanchuan, Hubei province, China. The UAV platform used was a Leica Aibot X6, featuring a 17-mm focal length Headwall Nano-Hyperspec sensor, operating at an altitude of 250 meters. The dataset comprises images with a size of 1217 × 303 pixels and a spatial resolution of approximately 0.109 meters, spanning 274 spectral bands from 400 to 1000 nm. The study area, a transitional zone between rural and urban land, includes buildings, water bodies, and fields with seven different crops: strawberry, cowpea, soybean, sorghum, water spinach, watermelon, and greens.
WHU-Hi-HongHu dataset [51] was collected on November 20, 2017 in Honghu City, Hubei province, China. The data acquisition was carried out using a DJI Matrice 600 Pro UAV equipped with a 17-mm focal length Headwall Nano-Hyperspec sensor. The study area is agriculturally diverse, featuring multiple crop types and various cultivars of the same crop, such as Chinese cabbage and cabbage, as well as different varieties of Brassica chinensis. The UAV operated at a flight altitude of 100 meters, capturing imagery with dimensions of 940 × 475 pixels, across 270 spectral bands ranging from 400 to 1000 nm. The spatial resolution of the images is approximately 0.043 meters.
IV-B Experiment Design
The experimental setup encompasses two comparative segments for classification evaluation. The first part involves basic model comparison experiments, while the second part focuses on comparing various tuning strategies.
In the basic model comparison part, we partition each target HSI dataset into two subsets: the training set and test set. We randomly extract 150 samples from each category as training samples and take the rest as the test samples in the Pavia University, Houston University and WHU-Hi datasets. For the Indian Pines dataset, we sample a small number of pixels from classes with scarce samples to ensure a more balanced representation for the experiments. Concretely, we randomly select 10 samples from each of the following six classes: Alfalfa (class 1), Grass/trees (class 5), Grass/pasture-mowed (class 7), Oats (class 9), Buildings-grass-trees (class 15), and Stone-steel towers (class 16). For the remaining classes, we sample 150 pixels from each class. In this Section IV.D, we compare our proposed Tri-Former model with other state-of-the-art algorithms.
In addition, we compare the proposed AGT approach with other popular tuning methods on HSIs collected from different sensors. The tuning strategies comparison results are presented in Section IV.G and Section IV.H. These experiments involve data from two source sensors: Salinas (captured by AVIRIS) and Pavia Center (captured by ROSIS). For the model trained on the data collected by one sensor, we fine-tune it on the target data collected by another sensor with a limited number of samples. We design different experimental groups, each containing 25, 50, and 75 labeled pixels per class, respectively. Furthermore, in Section IV.I, we evaluate the proposed method by cross-modality fine tuning experiments. Specifically, the basic Tri-Former model pretrained on RGB images is fintuned on different HSI datasets with different tuning method.
@@ IV-B Experiment Design
Figure 7: Results on Indian Pines. (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) MHCFormer; (i) MASSFormer; (j) Tri-Former.
IV-C Implementation Details
The optimization of the proposed method consists of two stages. The entire learning process is conducted on the server with 4 NVIDIA 3090 GPUs. During training, following the experience of previous works such as 3D-CNN, LWNet, etc. the training samples are cropped as $27\times 27$ spatial resolution patches, with a batch size of 96 on all three training dataset. We utilize the AdamW optimizer with a learning rate and weight decay value of $1\times 10^{-3}$ and $1\times 10^{-5}$, respectively. Following the training strategy proposed in ConvNext, the learning rate decays from $1\times 10^{-3}$ to $1\times 10^{-6}$ following the cosine scheduler. The warming-up stage comprises 5 epochs and all the models are trained for 300 epochs. To further tune the final network, we crop patches with spatial resolutions of $13\times 13$ and $27\times 27$, which are fed into the auxiliary branch and basic model, respectively. Batch sizes are set to 12 for all datasets during tuning.
TABLE III: Compared Methods Overview
Method 3D-CNN [^2] https://github.com/eecn/Hyperspectral-Classification [9] LWNet [^3] https://github.com/hkzhang91/LWNet [15] SpectralFormer [^4] https://github.com/danfenghong/IEEE_TGRS_SpectralFormer [11] SSFTT [^5] https://github.com/zgr6010/HSI_SSFTT [12] GraphGST [^6] https://github.com/yuanchaosu/TGRS-graphGST [14] MHCFormer [^7] https://github.com/Tikiten/MHCFormer [52] MassFormer [^8] https://github.com/hz63/MASSFormer [53]
Year TGRS16 TGRS19 TGRS21 TGRS22 TGRS24 TIE24 TGRS24
Architecture ConvNet ConvNet Transformer Transformer Transformer Transformer Transformer
@@ IV-C Implementation Details
TABLE IV: Comparison Experimental Results on Indian Pines Using 150 Training Samples Each Class
@@
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT MHCFormer MASSFormer Tri-Former
1 91.67 97.22 77.78 100.00 100.00 100.00 100.00 97.14
2 94.84 96.01 90.06 94.91 94.99 94.60 96.09 99.45
3 99.12 98.82 94.26 99.26 99.12 98.97 99.56 99.71
4 100.00 100.00 98.85 74.89 100.00 100.00 100.00 100.00
5 100.00 100.00 99.10 96.99 100.00 57.08 78.86 91.45
6 100.00 99.83 98.97 97.76 99.83 100 99.83 100.00
7 100.00 100.00 94.44 94.44 100.00 100.00 100.00 100.00
8 100.00 100.00 100.00 100.00 100.00 100.00 99.70 100.00
9 100.00 100.00 100.00 100.00 100.00 99.15 100.00 100.00
10 99.39 100.00 96.35 99.15 99.51 96.79 99.76 99.51
11 92.41 98.39 90.54 93.02 95.05 98.87 99.74 98.13
12 98.42 99.77 96.61 96.61 99.55 100 99.32 99.56
13 100.00 100.00 100.00 97.95 100.00 100.00 100.00 100.00
14 98.92 99.91 98.92 99.46 98.82 100 99.64 100.00
15 66.49 70.74 68.35 100.00 89.63 90.69 88.56 91.76
16 90.36 91.57 84.33 97.59 83.13 96.39 92.77 89.16
OA 95.23 97.46 93.08 96.02 97.12 95.36 97.48 98.41
AA 95.73 97.02 93.04 96.38 97.54 95.78 97.11 97.87
K 94.47 97.05 91.97 95.41 96.66 94.63 97.07 98.17
@@
TABLE V: Comparison Experimental Results on Pavia University Using 150 Training Samples Each Class
@@
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT MHCFormer MASSFormer Tri-Former
1 91.09 96.06 83.13 98.06 94.69 96.48 99.24 99.98
2 92.05 95.69 92.83 98.03 98.74 98.90 99.76 99.91
3 92.45 95.95 89.34 92.82 98.50 98.97 100.00 100.00
4 98.55 98.68 97.60 99.11 95.05 92.28 99.07 99.18
5 99.52 99.92 100.00 100.00 97.43 99.58 100.00 100.00
6 93.14 98.09 90.32 90.92 99.96 99.51 100.00 100.00
7 97.48 99.76 96.99 100.00 100.00 100.00 100.00 99.92
8 95.20 99.11 92.63 95.95 99.08 97.11 99.97 99.77
9 97.28 98.82 99.65 99.74 99.88 98.49 100.00 100.00
OA 93.26 96.86 91.65 98.06 98.03 98.02 99.71 99.88
AA 95.20 98.01 93.61 98.18 98.15 97.93 99.78 99.86
K 91.13 95.84 89.00 97.41 97.38 97.36 99.61 99.83
IV-D Comparison with state-of-the-art methods
In this section, we conduct a comparative analysis of our proposed Tri-Former against two ConvNet-based and five transformer-based HSI classification methods. The detail information about comparison methods are lists in Table III. Among them, ConvNet-based methods (e.g., 3D-CNN and LWNet) use fixed convolutional filters to extract spectral–spatial features but lack the adaptive long-range dependency modeling provided by transformers. The remaining methods are Transformer-based methods. Although these methods incorporate self-attention, they do not fully leverage the unique characteristics of HSIs. However, Our Tri-Former adopts a triplet structure (spectral, spatial, and 3D convolution branches) and uses an attention-gated tuning strategy (AGT) to bridge the domain gap, allowing cross-modal (RGB) pretraining.
The experimental results are presented in Table IV to Table IX, and the visual comparisons are shown in Fig. 7 to Fig. 12.
@@ IV-D Comparison with state-of-the-art methods
Figure 8: Results on Paiva University. (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) MHCFormer; (i) MASSFormer; (j) Tri-Former.
The quantitative analysis results of the aforementioned approaches on the Indian Pines dataset are listed in Table IV. Table IV shows that the proposed Tri-Former achieves the highest OA of 98.41%, outperforming other ConvNet-based HSI classification methods like 3D-CNN (95.23%), LWNet (97.46%), SpectralFormer (93.08%), GraphGST (96.02%) and SSFTT (97.12%). Among them, SSFTT incorporation of convolutional operations within the transformer structure, which benefits its performance. This improvement further substantiates some of the points we have mentioned in this work. Introducing convolutional operations to construct hybrid transformer models for processing HSI is a more suitable approach. The proposed Tri-Former further improved the OA metric of SSFTT by 1.29%, demonstrating its remarkable capability to accurately classify HSIs in this challenging dataset.
@@
TABLE VI: Comparison Experimental Results on Houston University Using 150 Training Samples Each Class
@@
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT MHCFormer MASSFormer Tri-Former
1 96.64 98.64 96.00 78.47 98.00 100.00 100.00 99.64
2 95.20 99.01 89.31 83.51 97.10 95.47 99.73 97.92
3 99.82 100.00 99.09 84.82 99.09 100.00 100.00 100.00
4 97.62 99.54 86.47 79.43 97.90 97.90 99.91 99.73
5 99.45 100.00 95.88 97.43 100.00 100.00 100.00 99.91
6 100.00 100.00 86.86 88.57 100.00 100.00 100.00 100.00
7 91.68 96.06 73.23 62.88 98.66 95.01 98.93 99.11
8 94.70 96.71 85.56 54.39 97.62 96.07 100.00 97.17
9 96.82 97.19 85.57 54.38 99.82 96.82 99.36 98.82
10 98.33 99.44 91.55 70.10 100.00 99.81 99.91 99.63
11 97.05 100.00 92.35 51.34 100.00 100.00 99.72 97.05
12 96.12 98.25 86.98 60.85 96.21 98.43 99.63 100.00
13 95.61 100.00 82.45 86.20 99.37 100.00 100.00 100.00
14 100.00 100.00 100.00 94.96 100.00 100.00 100.00 100.00
15 100.00 100.00 97.65 86.67 100.00 99.8 100.00 100.00
OA 96.58 98.70 89.19 94.42 98.69 98.23 99.76 99.05
AA 97.14 98.99 89.93 95.13 98.92 98.63 99.81 99.26
K 96.29 98.59 88.28 93.94 98.57 98.08 99.74 98.97
TABLE VII: Comparison Experimental Results on WHU-Hi-Longkou Using 150 Training Samples Each Class
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT MHCFormer MASSFormer Tri-Former
1 99.39 97.28 98.97 99.87 96.21 99.47 99.67 99.74
2 98.39 98.12 86.02 99.67 95.39 99.27 99.98 99.87
3 99.86 100.00 99.27 100.00 99.44 99.60 100.00 100.00
4 94.22 93.80 81.32 98.62 94.00 97.42 98.78 98.26
5 100.00 100.00 98.17 99.80 99.65 100.00 99.85 99.38
6 99.39 97.56 98.73 98.09 99.03 99.22 99.69 100.00
7 99.82 99.20 99.05 97.87 99.94 98.39 98.98 99.81
8 96.95 93.92 94.83 95.97 96.62 92.13 94.45 96.53
9 98.44 98.82 98.91 96.20 96.24 97.78 98.48 98.76
OA 97.80 96.90 92.84 98.48 97.01 98.17 98.98 99.18
AA 98.50 97.63 95.03 98.48 97.39 98.14 98.88 99.15
K 97.12 95.94 90.80 98.01 96.09 97.61 98.66 98.93
Figure 9: Results on Houston University. (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) MHCFormer; (i) MASSFormer; (j) Tri-Former.
TABLE VIII: Comparison Experimental Results on WHU-Hi-Hanchuan Using 150 Training Samples Each Class
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT MHCFormer MASSFormer Tri-Former
1 92.21 89.85 87.06 98.32 93.14 97.21 98.55 97.82
2 95.84 95.09 84.90 95.34 82.14 92.92 98.35 98.13
3 93.01 93.94 91.71 99.96 94.83 98.92 99.99 99.72
4 96.75 95.87 96.90 99.85 97.71 99.44 99.65 99.62
5 100.00 100.00 100.00 100.00 99.43 100.00 100.00 100.00
6 90.12 87.66 84.95 98.29 91.95 93.63 99.61 97.26
7 96.96 96.44 96.68 99.84 96.84 98.73 99.65 99.91
8 90.50 85.05 81.43 95.67 86.37 93.99 95.87 98.60
9 92.64 86.40 91.50 95.45 93.23 90.02 97.69 97.86
10 92.92 89.92 96.69 97.69 99.22 98.76 99.23 99.44
11 96.92 94.31 95.63 97.03 95.87 97.70 96.97 99.45
12 98.53 97.00 99.18 100.00 96.32 99.86 100.00 99.83
13 89.46 85.09 88.04 92.17 82.52 84.72 92.65 95.61
14 96.95 94.20 87.54 97.61 93.26 95.06 97.27 97.72
15 100.00 98.07 99.47 100.00 93.81 100.00 99.90 100.00
16 95.70 96.28 98.09 98.98 99.09 98.04 99.48 99.07
OA 94.38 92.76 91.76 97.75 93.81 96.50 98.37 98.56
AA 94.91 92.82 92.49 97.89 93.48 96.24 98.43 98.75
K 93.45 91.56 90.41 97.37 92.76 95.93 98.09 98.31
TABLE IX: Comparison Experimental Results on WHU-Hi-Honghu Using 150 Training Samples Each Class
Models 3D-CNN LWNet SpectralFormer GraphGST SSFTT MHCFormer MASSFormer Tri-Former
1 94.46 91.75 95.98 99.05 98.79 96.87 96.31 97.34
2 97.29 97.95 94.68 98.13 95.54 95.12 97.44 98.30
3 90.55 91.82 90.55 95.60 92.11 94.30 92.53 95.99
4 96.71 95.02 96.09 99.03 96.15 98.97 97.50 99.30
5 98.52 98.38 95.59 100.00 97.82 99.77 99.70 99.51
6 94.77 91.94 91.24 98.95 96.52 98.51 97.11 98.62
7 79.01 70.19 77.25 95.02 87.22 87.66 91.38 93.04
8 92.90 96.62 76.57 100.00 96.41 99.72 99.23 100.00
9 97.82 94.26 98.76 98.97 98.65 98.76 96.28 99.22
10 94.79 94.89 80.35 98.14 88.36 98.50 98.46 99.26
11 92.31 83.66 80.49 96.30 93.33 96.70 96.02 96.84
12 90.57 85.81 83.72 98.86 85.29 95.35 96.71 97.40
13 91.39 83.51 77.12 95.06 87.41 93.62 98.08 96.77
14 96.14 97.21 89.52 99.33 91.35 97.63 97.00 99.18
15 100.00 99.41 98.60 99.88 99.30 99.88 100.00 100.00
16 99.76 99.25 95.92 99.80 94.05 99.65 98.02 100.00
17 98.85 98.64 97.64 99.79 98.50 100.00 99.72 99.51
18 98.79 97.75 96.58 99.05 99.05 98.89 98.04 98.73
19 92.78 89.22 92.72 98.91 95.59 96.98 98.49 98.91
20 97.54 99.70 97.79 99.22 97.99 99.79 98.65 99.10
21 99.75 99.92 98.04 100.00 98.47 100.00 100.00 100.00
22 98.43 99.41 96.73 100.00 98.77 100.00 99.25 99.77
OA 94.37 91.83 91.40 98.29 94.46 97.35 96.44 98.31
AA 95.14 93.47 91.00 98.59 94.85 97.58 97.29 98.49
K 92.92 89.75 89.24 97.84 93.04 96.65 95.51 97.87
Figure 10: Results on WHU-Hi-LongKou. (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) MHCFormer; (i) MASSFormer; (j) Tri-Former.
Fig. 7 shows the corresponding visual results. At the top, a color-composite image is presented, accompanied by color codes that represent different types of ground objects. The bottom-right corner features a magnified view, aimed at providing a clearer comparison of the classification performance of different methods. From Fig. 7, we can see that all methods exhibit varying degrees of misclassification. Our proposed Tri-Former demonstrates the least amount of misclassification, approaching the ground truth closely.
Figure 11: Results on WHU-Hi-HanChuan . (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) MHCFormer; (i) MASSFormer; (j) Tri-Former.
Figure 12: Results on WHU-Hi-HongHu. ( (a) Composite color map; (b) Ground truth; (c) 3D-CNN; (d) LWNet; (e) SpectralFormer; (f) GraphGST; (g) SSFTT; (h) MHCFormer; (i) MASSFormer; (j) Tri-Former.
Similar trends are also demonstrated in the experiments conducted on the Pavia University and Houston University datasets. When evaluated on the Pavia University dataset, the proposed Tri-Former algorithm surpasses the state-of-the-art method SSFTT, achieving remarkable improvements in key evaluation metrics. Specifically, we observe substantial enhancements in Overall Accuracy (OA), Average Accuracy (AA), and Kappa (K) coefficient by 1.85%, 1.71%, and 2.45%, respectively. On the Houston University dataset, where the potential for performance improvement is considerably limited, our Tri-Former still manages to achieve the best performance among all methods compared. Fig. 8 and Fig. 9 are the corresponding visual comparison results. On these two datasets, especially on the Pavia University dataset, the proposed Tri-Former produces classification result maps that closely resemble the groundtruth.
Across all three WHU-Hi datasets, the proposed Tri-Former consistently outperforms conventional models like 3D-CNN, LWNet, and even newer transformer-based models such as SpectralFormer, GraphGST, and SSFTT. This can be attributed to Tri-Former’s ability to effectively capture both spatial and spectral features while handling the unique challenges of hyperspectral imaging data. The results from Table VII to Table IX substantiate the argument that incorporating hybrid transformer structures, as demonstrated in Tri-Former, is a highly effective approach for HSI classification.
IV-E Inference speed analysis on Tri-Former